# Hal-Eval: A Universal and Fine-grained Hallucination Evaluation Framework for Large Vision Language Models

### v1

@@ 3.2.2 Construction of Hal-Data
This dataset named Hal-Data which is generated based on the automatic hallucination annotation pipline AFHA and comprises 2 million instances. Hal-Data comprises two parts: 130K GPT-4 annotated and curated image-text pairs named Hal-Data 130k, with each pair consisting of an image, a valid image caption, and a description containing hallucinations. The second part consists of 2M image-text pairs named Hal-Data 2M, crafted by our caption model trained on the 130K high-quality captions. In the following section, we provide a detailed exposition of the process involved in creating the Hal-Data.
@@
Data Collection for Hal-Data 130k To maximize the diversity and comprehensiveness of our data, we initially compiled approximately 200K images from various sources, including 80K images from the in-domain COCO dataset Lin et al. (2014) and 80K web images, such as those from CC Changpinyo et al. (2021), SBU Ordonez et al. (2011), and LAION Schuhmann et al. (2022). Additionally, to better align with the style of LVLM outputs, we also collected 40K image-text datasets from ShareGPT4-V Chen et al. (2023b). Subsequently, we employed AFHA to annotate this portion of the data, resulting in a final collection of 130K GPT4 meticulously annotated instances and named it as Hal-Data 130k.
@@
Generation for Hal-Data 2M. Building upon the Hal-Data 130k dataset, we endeavored to further expand the scale of our dataset. Due to the high cost associated with using GPT-4, we leveraged the Hal-Data 130k dataset to fine-tune the currently open-source large-scale language model LLaMA2 13B Touvron et al. (2023a), resulting in a hallucination data annotation model named Hal-Annotator. Thanks to its training on diverse and comprehensive data, the Hal-Annotator is capable of generating highly quality and content-related annotations. This approach allows the data scaling phase to proceed without the need for using the paid GPT-4. To accumulate a substantial volume of high-quality image-text pairs, we selected a subset of 2 million image-caption pairs from current public datasets (see Appendix A.1 for more details) and employed our pre-trained Hal-Annotator to modify the image captions by introducing different types of hallucinations and annotating them.
3.2.3 Instruction tuning of Hal-Evaluator.
[…]
3.2.4 Evaluation Metric
@@
To evaluate the generative hallucination of LVLMs, we prompt them to describe images from both an in-domain 5K dataset and an out-of-domain 5K dataset mentioned in Subsection 3.1.1 with short length and longer length. These descriptions, coupled with the respective images, are then fed into the pre-trained Hal-Evaluator. Our procedure involves prompting Hal-Evaluator to evaluate the existence and category of any hallucinatory content. Accuracy serves as the principal metric for our evaluation, which measures the proportion of outputs correctly identified as free from hallucinations. Suppose the number of all outputs is $N$, the outputs that contains hallucinations is $N_{h}$, the accuracy is calculated as $A=\frac{N-N_{h}}{N}$. Moreover, we track the probability of various types of hallucinations, encapsulated in the hallucination ratio. For instances, the number of output contains the object hallucination is $N^{o}_{h}$, the object ratio $r_{o}$ is calculated as $r_{O}=\frac{N^{o}_{h}}{N_{h}}$.
4 Experiments
Hal-Eval is divided into two distinct segments: Discriminative Evaluation and Generative Evaluation. We have opted to assess five widely utilized open-source LVLMs: MiniGPT-4 Zhu et al. (2023), InstructBLIP Dai et al. (2023), mPLUG-owl Ye et al. (2023), LLaVA Liu et al. (2023e), LLaVA1.5 Liu et al. (2023c).
@@ 4 Experiments
LLaVA 1.5 57.7 77.2 21.9 34.2 14.2
Table 1: Results of LVLMs under evaluation of four hallucination types on the in-domain dataset and out-of-domain dataset. Yes denotes the proportion of answering “Yes” to the given question.
@@
Table 2: Generative Hallucination Evaluation for LVLMs.
4.1 Discriminative Evaluation
4.1.1 Main Results
@@
As shown in the Table 1, we evaluate the performance of five models, on different types of hallucinations, following the method outlined in Subsection 3. First, we noticed that the results derived from the POPE Li et al. (2023e) metric indicate a significant tendency among most models to favor "yes" responses. In contrast, within our discriminative evaluations, such a penchant is exclusively noted in the InstructBLIP model. This distinction serves to underscore the efficiency of our evaluation strategy in attenuating the inclination towards "yes" answers that is frequently encountered in LVLMs. Moreover, LLaVA1.5 and LLaVA exhibit a more pronounced predilection for hallucinations when tested against out-of-domain datasets as opposed to in-domain datasets. This trend could possibly be ascribed to the prevalent incorporation of in-domain imagery during the instruction tuning phase of the models.
4.1.2 Analysis of Discriminative Evaluation
@@
Data Reliability Analysis. Our proposed evaluation dataset comprises 5,000 in-domain and 5,000 out-of-domain images, which we annotated based on the AFHA framework. To verify the accuracy of the annotations, we randomly sampled 100 cases each from both the in-domain and out-of-domain data for manual validation. As shown in the table, we found that after GPT-4’s annotation and filtering process, the annotation accuracy rate in the in-domain dataset reached 98%. Meanwhile, the annotation accuracy for the out-of-domain dataset stood at 97%. This high level of accuracy in both datasets underscores the effectiveness of our annotation process.
Figure 4: Comparison of LLaVA1.5 and LLaVA 1.5-COT. We report the F1 score for both of them.
@@
Effectiveness of Chain-Of-Thought (COT) for Mitigating Discriminative Hallucination. For discriminative evaluation, we employed a chain of thought (COT) approach to systematically evaluate whether the LVLM matches the content of images with their respective captions (Refer to Appendix B.4 for more details). As shown in the Figure 4, we observed a significant reduction in discriminative hallucinations on both in-domain and out-of-domain datasets after employing COT to LLaVA 1.5. Our experiments indicate that utilizing COT is particularly effective in reducing discriminative hallucinations for LVLMs, especially for those related to relationships and events. We posit that the complexity inherent in relationship and event type hallucinations arises from their intrinsic reliance on contextual and inferential understanding.
4.2 Generative Evaluation
4.2.1 Main Results
@@
As indicated in Table 2, our investigation has revealed that contemporary models continue to have a significant inclination toward producing hallucinations. MiniGPT-4 and InstructBLIP, displayed robust in-domain accuracy, with the latter achieving 89.1% accuracy when the average output length was approximately 10 tokens. Both mPLUG-owl and LLaVA showed moderate performance across all evaluated metrics, whether tested on in-domain or out-of-domain data. Furthermore, we found that (1) When models generate shorter outputs, the majority of hallucination content produced by the models predominantly consisted of objects. (2).As the length of the output increases, all models became more prone to producing hallucination content, with the incidence of event hallucinations rising markedly.
@@
Table 3: Correlation between each evaluation metric and human judgment on LVLM hallucinations, measured by Pearson’s $r$, Spearman’s $\rho$, and Kendall’s $\tau$.
@@
Figure 5: The left sub-figure displays the results of the discriminative evaluation for GPT-4V and Hal-Evaluator. The right sub-figure compares the ROUGE-L between hallucination content detected by GPT-4V and Hal-Evaluator with the annotated hallucination content.
@@
Figure 6: The figure depicts the proportions of different types of hallucinations in the outputs of InstructBLIP and LLaMA1.5, as well as the gray line illustrating the variation in accuracy.
4.2.2 Analysis of Generative Evaluation
@@
Correlations with Human Evaluations. To verify the correlation between the generative evaluation and human judgment, we following experiments. We firstly select the test dataset from COCO 2014 Lin et al. (2014) for human evaluation. This test set comprises 50 images. Each image is supplemented with five reference captions and object labels provided by the COCO dataset. We selected three LVLMs – LLaVA Liu et al. (2023e), mPLUG-owl Ye et al. (2023), and instructBLIP Dai et al. (2023) – to describe the content of the test set, and we sought the annotation and evaluation from 15 human annotators to evaluate the presence of hallucinations in these data. We made a comparison among five benchmarks: ROUGE-L Lin (2004), BLEU-4 Papineni et al. (2002), CHAIR Rohrbach et al. (2018), Hal-EML Wang et al. (2023b), GPT4-V along with the module of Hal-Eval – Hal-Eval-Generative. Table 3 delineates the correlation between various evaluation metrics and human judgment regarding LVLM faithfulness, gauged using Pearson’s $r$, Spearman’s $\rho$, and Kendall’s $\tau$. Among the diverse array of metrics, our generative evalutioin metric distinctly stand out which demonstrate a robust positive correlation, underlining the superior alignment with human perceptions.
@@
Effectiveness of Hal-Evaluator. To further verify the effectiveness of Hal-Evaluator for hallucination detection, we evaluated Hal-Evaluator and GPT-4V (as candidate LVLMs instead of evaluator here) based on the discriminative approach scripted in subsection 3.1.2, evaluating the detection of different types of hallucinations in 5K in-domain data. The results disclosed that Hal-Evaluator outperforms GPT-4V in hallucination detection ability, as shown in the Sub-Figure 5 (a).
Analysis of Event Hallucination. We tasked GPT-4V and Hal-Evaluator with identifying hallucination content in image descriptions of in-domain evaluation dataset of Hal-Eval. (Noted that our evaluation dataset was also carefully annotated to mark hallucination content within the descriptions.) We evaluated the overlap between the hallucination content as identified by GPT-4V and Hal-Evaluator and the actual hallucination information that exists in the dataset containing hallucinations. The overlap was quantified using the ROUGE-L score, as shown in Sub-figure 5 (b). The experiment result indicates that, for the first three types of hallucinations (object, attribute, relation), both GPT-4V and Hal-Evaluator are capable of accurately identifying the majority of hallucination content from image descriptions. However, for event hallucinations, GPT-4V fails to precisely locate the hallucination content, whereas Hal-Evaluator exhibits an accurate identifying ability. This discrepancy underscores the complexity inherent to event hallucinations and concurrently attests to the reliability of Hal-Evaluator. Additionally, Hal-Evaluator possesses enhanced evaluation capabilities, allowing it to reduce reliance on external APIs such as GPT-4V. Furthermore, Hal-Evaluator also can mitigate hallucinations in LVLMs by correcting hallucination content (See Appendix C.3).
Impact of LVLM output length on Generative Hallucination. In this section, we investigate the potential correlation between the output length of LVLMs and the occurrence of generative hallucinations. We conducted the following experiments on LLaVA 1.5 and InstructBLIP. Experiments were carried out on LLaVA 1.5 and InstructBLIP wherein both models were prompted to describe the content of images with outputs of varying token lengths, using the Hal-Evaluator to detect and analyze the outputs at each length. The proportions of different types of hallucinations and the accuracy of outputs without hallucinations are visualized in the Figure 6. The following findings were observed: (1) The probability of hallucinations occurring in brief responses is generally lower. (2) With the increase in output length, the occurrence of event hallucinations significantly rises. These results shed light on the relationship between output length and the incidence of event hallucinations.
4.3 Utilizing Hal-Data for Supervised Fine-tuning
Figure 7: Radar Chart of Different Models on Multiple General Benchmarks
@@ 4.3 Utilizing Hal-Data for Supervised Fine-tuning
in-domain out-of-domain in-domain out-of-domain length
@@
Table 4: The evaluation results of Hal-VL on Hal-Eval. For the discriminative evaluation, we have only listed the average F1 scores for different types of hallucination. For the generative evaluation, we list the accuracy with short output length.

### v2

@@ 4.3.2. Instruction finetuning of Hal-Evaluator
This dataset, referred to as Hal-Data, was generated using an automatic hallucination annotation pipeline and comprises 2 million instances. Hal-Data is split into two parts: Hal-Data 130k, which includes 130,000 GPT-4 annotated and curated image-text pairs, each consisting of an image, a valid image caption, and a hallucination description; and Hal-Data 2M, which includes 2 million image-text pairs created by our caption model trained on the 130,000 high-quality captions from Hal-Data 130k. Below, we detail the creation process for Hal-Data.
@@
Data Collection for Hal-Data 130k: To ensure diversity and comprehensiveness, we initially compiled about 200,000 images from various sources, including 80,000 in-domain COCO dataset images (Lin et al., 2014), 80,000 web images from sources like CC (Changpinyo et al., 2021), SBU (Ordonez et al., 2011), and LAION (Schuhmann et al., 2022), and 40,000 image-text datasets from ShareGPT4-V (Chen et al., 2023a) to match the style of LVLM outputs. We then used AFHA to annotate this data, resulting in a final collection of 130,000 meticulously annotated GPT4 instances, named Hal-Data 130k.
@@
Generation for Hal-Data 2M: We further selected a subset of 2 million image-caption pairs from current public datasets (see Appendix B.2 for more details) and constructed a large-scale hallucination dataset named Hal-Data 2M. Due to the high cost of using GPT-4, we fine-tuned the open-source large-scale language model LLaMA2 13B (Touvron et al., 2023a) on Hal-Data 130k and employed it to modify the image captions of Hal-Data 2M by introducing different types of hallucinations and annotating them.
Based on Hal-Data, we fine-tuned LLaVA 1.5 13B (Liu et al., 2023c), recent SOTA LVLM, with 2M instruction data specifically designed for detecting and correcting hallucinations in image captions, leading to the development of Hal-Evaluator. (For more details, please refer to our supplemental material D.1.)
4.3.3. Generative Evaluation Based in Hal-Evaluator
As illustrated in Figure 3, the input to Hal-Evaluator consists of two parts: an image and the corresponding textual description by the candidate LVLM to be evaluated. We prompt Hal-Evaluator to first determine if the text description contains hallucinations based on the image content. If hallucinations are detected, Hal-Evaluator will identify the type of hallucination and its content. Ultimately, Hal-Evaluator can also correct the hallucinatory content in the image description, providing a revised depiction of the image.
4.3.4. Evaluation Metric
@@
To evaluate the generative hallucination of LVLMs, we prompt them to describe images from both an in-domain 5K dataset and an out-of-domain 5K dataset mentioned in Subsection 4.2.1 with short length and longer length. These descriptions, coupled with the respective images, are then fed into the pre-trained Hal-Evaluator. Our procedure involves prompting Hal-Evaluator to evaluate the existence and category of any hallucinatory content. Accuracy serves as the principal metric for our evaluation, which measures the proportion of outputs correctly identified as free from hallucinations. Suppose the number of all outputs is $N$, the outputs that contain hallucinations are $N_{h}$, and the accuracy is calculated as $A=\frac{N-N_{h}}{N}$. Moreover, we track the probability of various types of hallucinations encapsulated in the hallucination ratio. For instance, the number of outputs containing the object hallucination is $N^{o}_{h}$, and the object ratio $r_{o}$ is calculated as $r_{O}=\frac{N^{o}_{h}}{N_{h}}$.
@@
LLaVA 1.5 57.7 77.2 41.9 44.2 14.2
GPT4-V 85.3 83.2 84.5 84.5 50.3
Table 1. Results of LVLMs under evaluation of four hallucination types on the in-domain dataset and out-of-domain dataset. Yes denotes the proportion of answering “Yes” to the given question.
5. Experiments
Hal-Eval is divided into two distinct segments: Discriminative Evaluation and Generative Evaluation. We have opted to assess five widely utilized open-source LVLMs: MiniGPT-4 (Zhu et al., 2023a), InstructBLIP (Dai et al., 2023), mPLUG-owl (Ye et al., 2023), LLaVA (Liu et al., 2023e), LLaVA1.5 (Liu et al., 2023c) and one close-source LVLM: GPT4-V (OpenAI, 2023b).
@@ 5. Experiments
[…]
80.2 32.9 21.0 16.6 30.4 77.6 30.9 18.0 13.6 38.4 73.1
Table 2. Generative Hallucination Evaluation for LVLMs.
5.1. Discriminative Evaluation
5.1.1. Main Results
@@
As shown in Table 1, we evaluate the performance of five models on different types of hallucinations following the method outlined in Subsection 3. First, LLaVA1.5 and LLaVA exhibit a more pronounced predilection for hallucinations when tested against out-of-domain datasets as opposed to in-domain datasets. This trend could possibly be ascribed to the prevalent incorporation of COCO (Lin et al., 2014) during the instruction tuning phase of the models. Moreover, we noticed that the results derived from the POPE (Li et al., 2023a) metric indicate a significant tendency among most models to favor ”yes” responses. In contrast, within our discriminative evaluations, such a penchant is exclusively noted in the InstructBLIP. This distinction serves to underscore that Hal-eval can effectively avoid the bias of the model towards answering ”yes”. In the end, our findings indicate that, with the exception of GPT-4, the performance of the currently accessible open-source LVLMs in discriminative evaluations is subpar. These models face challenges in accurately discerning and interpreting hallucinatory content within image descriptions.
5.1.2. Analysis of Discriminative Evaluation
@@
Data Reliability Analysis: Our evaluation dataset comprises 5,000 in-domain and 5,000 out-of-domain images, which we annotated based on the AFHA framework. To verify the accuracy of the annotations, we randomly sampled 100 cases each from both the in-domain and out-of-domain data for manual validation (Please refer to supplemental material C for more details). We found that after GPT-4’s annotation and filtering process, the annotation accuracy rate in the COCO dataset reached 98%. Meanwhile, the annotation accuracy for the out-of-domain dataset stood at 97%. This high level of accuracy in both datasets underscores the effectiveness of our annotation process.
Figure 4. Comparison of LLaVA1.5 and LLaVA 1.5-COT. We report the F1 score for both of them.
@@
Effectiveness of Chain-Of-Thought (COT) for Mitigating Discriminative Hallucination: For discriminative evaluation, we employed a chain of thought (COT) approach to evaluate whether the LVLM matches the content of images with their respective captions (Appendix C.4). As shown in Figure 4, we observed a significant reduction in discriminative hallucinations on both in-domain and out-of-domain datasets after employing COT to LLaVA 1.5. Based on the above experiments, we have the following finding:
Finding 1: Utilizing COT is particularly effective in reducing discriminative hallucinations for LVLMs, especially for those related to relationships and events.
We suggest that the increased effectiveness of COT with relationship and event type hallucinations is due to their intrinsic reliance on contextual understanding and inference-making.
5.2. Generative Evaluation
5.2.1. Main Results
@@
As indicated in Table 2, our investigation has revealed that contemporary models continue to have a significant inclination toward producing hallucinations. MiniGPT-4 and InstructBLIP, displayed robust in-domain accuracy, with the latter achieving 89.1% accuracy when the average output length was approximately 10 tokens. Both mPLUG-owl and LLaVA showed moderate performance across all evaluated metrics, whether tested on in-domain or out-of-domain data. GPT-4V achieved the best result, but we also observed a notable decline in accuracy with increasing output length, accompanied by a significant rise in the proportion of event-type hallucinations. Furthermore, we found that when generating long responses, all models became more prone to producing hallucination content, with the incidence of event hallucinations rising markedly.
@@
Table 3. Correlation between each evaluation metric and human judgment on LVLM hallucinations, measured by Pearson’s $r$, Spearman’s $\rho$, and Kendall’s $\tau$.
@@
Figure 5. The left sub-figure displays the results of the discriminative evaluation for GPT-4V and Hal-Evaluator. The right sub-figure compares the ROUGE-L between hallucination content detected by GPT-4V and Hal-Evaluator with the annotated hallucination content.
@@
Figure 6. The figure depicts the proportions of different types of hallucinations in the outputs of InstructBLIP and LLaMA1.5, as well as the gray line illustrating the variation in accuracy.
5.2.2. Analysis of Generative Evaluation
@@
Correlations with Human Evaluations: To verify the correlation between the generative evaluation and human judgment, we conduct the following experiments. We first select the test dataset from COCO 2014 (Lin et al., 2014) for human evaluation. This test set comprises 50 images. Each image is supplemented with five reference captions and object labels provided by the COCO dataset. We selected three LVLMs – LLaVA (Liu et al., 2023e), mPLUG-owl (Ye et al., 2023), and instructBLIP (Dai et al., 2023) – to describe the content of the test set, and we sought the annotation and evaluation from 15 human annotators to evaluate the presence of hallucinations in these data. We made a comparison among five benchmarks: ROUGE-L (Lin, 2004), BLEU-4 (Papineni et al., 2002), CHAIR (Rohrbach et al., 2018), Hal-EML (Wang et al., 2023c), GPT4-V along with the module of Hal-Eval – Hal-Eval-Generative. Table 3 delineates the correlation between various evaluation metrics and human judgment regarding LVLM faithfulness, gauged using Pearson’s $r$, Spearman’s $\rho$, and Kendall’s $\tau$. Our generative evaluation distinctly stands out and demonstrates a robust positive correlation, underlining the superior alignment with human perceptions.
@@
Effectiveness of Hal-Evaluator: To further verify the effectiveness of Hal-Evaluator for hallucination detection, we evaluated Hal-Evaluator and GPT-4V (as candidate LVLMs instead of evaluator here) based on the discriminative evaluation of Hal-Eval scripted in subsection 4.2.2, evaluating the detection of different types of hallucinations in 5K coco data. The results disclosed that Hal-Evaluator outperforms GPT-4V in hallucination detection ability, as shown in the Sub-Figure 5 (a).
Discriminative (Ave F1) Generative (Hallucination Ratio)
Object Relation Attribute Event Object Relation Attribute Event
52.8 52.6 51.8 54.3 38.3 12.7 8.8 41.2
Table 4. This table displays the average F1 scores for various hallucination types in discriminative assessments and the average hallucination rates for LVLM’s long outputs (>=>=> =80) in generative assessments.
@@
ID OOD ID OOD length
@@
Table 5. The evaluation results of Hal-VL on Hal-Eval. For the discriminative evaluation, we have only listed the average F1 scores for different types of hallucination. For the generative evaluation, we list the accuracy with short output length.
General VQA General VQA (Zero-shot)
Method #Params VQAv2 GQA VizWizQA TextVQA SciQA
InstructBLIP (Dai et al., 2023) 8.2B - 49.2 34.5 $50.1^{\dagger}$ 60.5
Shikra 7.2B 77.4 - - - -
Qwen-VL-Chat (Bai et al., 2023a) 9.6B 78.2 57.5 38.9 $61.5^{\ddagger}$ 68.2
LLaVA (Liu et al., 2023e) 7.2B 71.3 41.3 36.7 $50.2^{\dagger}$ 61.5
MiniGPT-4 (Zhu et al., 2023a) 7.2B 65.2 30.8 30.2 $52.3^{\dagger}$ 58.4
LLaVA1.5 (Liu et al., 2023c) 7.2B 78.5 62.0 50.0 $58.2^{\dagger}$ 66.8
[…]
Table 6. Performance comparison on visual question answering. For VQA, accuracy is reported. Note that specialists are fine-tuned on each individual dataset. † denotes OCR inputs are utilized. ‡ indicates the model has trained on the dataset.
Analysis of Event Hallucination: We tasked GPT-4V and Hal-Evaluator with detecting the hallucination content in image descriptions of the evaluation dataset of Hal-Eval. We evaluated the overlap between the hallucination content as identified by GPT-4V and Hal-Evaluator and the annotated hallucination content that exists in the descriptions. The overlap was quantified using the ROUGE-L score, as shown in Sub-figure 5 (b). The experimental results show that both GPT-4V and Hal-Evaluator can accurately identify the majority of hallucinated content in image descriptions for the first three types of hallucinations (object, attribute, relation). However, when it comes to event hallucinations, GPT-4V struggles to pinpoint the hallucinated content accurately, while Hal-Evaluator demonstrates a reliable identifying capability. Considering that GPT-4V is the current top-performing LVLM yet its difficulty in accurately identifying event hallucinations, this underscores the intrinsic complexity of event hallucinations and Hal-Evaluator’s reliability in detecting event-type hallucinations.
Impact of LVLM output length on Generative Hallucination. We investigate the potential correlation between the output length of LVLMs and the occurrence of generative hallucinations. We conducted the following experiments on LLaVA 1.5 and InstructBLIP. Experiments were carried out on LLaVA 1.5 and InstructBLIP wherein both models were prompted to describe the content of images with outputs of varying token lengths, using the Hal-Evaluator to detect and analyze the outputs at each length. The proportions of different types of hallucinations and the accuracy of outputs without hallucinations are visualized in Figure 6. Combining the results in subsection 5.2.1, the following findings were observed:
Finding 2: The incidence of hallucinations, especially event hallucinations, increases with the length of the output. Length control becomes a crucial aspect of generative evaluations, affecting comparative performance trends among LVLMs under varied output lengths.
5.3. Comparative Analysis of Discriminative and Generative Evaluations
As shown in Table 4, for event hallucination, most LVLMs perform better in discriminative evaluation compared to other hallucination types. For example, we calculated the average F1 scores for discriminative evaluations across all models for different types of hallucinations, and the average F1 score for event types was 54.3, which is higher than that for other hallucinations. This suggests that models are more resistant to event-type hallucinations. However, when evaluating the models using a generative evaluation approach, we observed a higher frequency of event-type hallucinations in longer output sequences, contradicting the results from discriminative evaluations. We believe this is mainly because event-type hallucinations often contain more complex and rich information inconsistent with the image content, making them easier for models to handle in discriminative evaluations. However, this does not accurately reflect the models’ ability to avoid event-type hallucinations effectively. Herefore, generative evaluation is a more effective method for assessing event-type hallucinations. For other types of hallucinations, such as objects, attributes, and relationships, research (Li et al., 2023a)(Lovenia et al., 2023)(Hu et al., 2023) indicates that discriminative evaluations are sufficiently effective in reflecting whether models tend to such hallucinations. In summary, we have the following findings:
Finding 3: The suitability of evaluation methodology varies according to the type of hallucinations. Using both discriminative and generative evaluations together gives a fuller view of tendencies to LVLM hallucination.
5.4. Utilizing Hal-Data for Supervised Fine-tuning